# Training Machine Learning models at the Edge: A Survey

### v2

@@ Abstract
Edge Computing (EC) has gained significant traction in recent years, promising enhanced efficiency by integrating Artificial Intelligence (AI) capabilities at the edge. While the focus has primarily been on the deployment and inference of Machine Learning (ML) models at the edge, the training aspect remains less explored. This survey delves into Edge Learning (EL), specifically the optimization of ML model training at the edge. The objective is to comprehensively explore diverse approaches and methodologies in EL, synthesize existing knowledge, identify challenges, and highlight future trends. Utilizing Scopus’ advanced search, relevant literature on EL was identified, revealing a concentration of research efforts in distributed learning methods, particularly Federated Learning (FL). This survey further provides a guideline for comparing techniques used to optimize ML for edge learning, along with an exploration of different frameworks, libraries, and simulation tools available for EL. In doing so, the paper contributes to a holistic understanding of the current landscape and future directions in the intersection of edge computing and machine learning, paving the way for informed comparisons between optimization methods and techniques designed for edge learning.
@@
Machine Learning; Edge Computing; Edge AI; Edge Learning; On-Device Training; Artificial Intelligence; IoT.
publicationid: pubid:
@@ I Introduction
In the past years, the fields of Artificial Intelligence (AI) and Machine Learning (ML) has witnessed significant growth and advancements, demonstrating remarkable success across various industrial applications [1]. The essence of ML has always been shaped by the interplay between data and algorithmic models. Large quantities of data being often required to effectively train ML models. And traditionally, the datasets were collected in cloud storage buckets, large databases and data lakes, and processed in central cloud servers, to train various machine learning and deep learning models.
@@
Conversely, The rapid proliferation of smart devices and sensors in recent years has led to an explosion of data generation at the edge of the network. With edge devices generating vast quantities of data outside the traditional cloud servers, the raising awareness of the privacy and security concerns that come with sending these data to the cloud, as well the desire to optimize the bandwidth consumption on the increasing number of edge devices and reduce the computational load on cloud servers, have precipitated a paradigm shift towards Edge Computing (EC). In this context, computational processes are decentralized and migrated to edge devices, setting the stage for a novel intersection between machine learning and edge computing.
@@
This paradigm shift has spurred an increasing interest towards Edge Machine Learning (Edge ML). A union between machine learning and edge computing, where ML models are deployed at the edge, closer to end devices, for both inference and training purposes. The latter, known as Edge Learning (EL), involves the training of ML models directly at the edge. Traditionally, ML models have relied heavily on centralized cloud infrastructure for both training and deployment. However, this approach poses several challenges, including high latency, significant communication overheads, and concerns around data privacy and security. By moving the computation closer to the source of the data, EL offers a promising solution to address these issues while enabling real-time decision-making and reducing the load on cloud resources. Furthermore, edge learning creates opportunities for innovative applications of ML, such as privacy aware recommendation, smart technologies, and autonomous vehicles, spanning across multiple industries from healthcare and manufacturing to the agriculture and space industry.
@@
However, training ML models at the edge presents unique challenges due to the limited computational capabilities and memory constraints of edge devices. Moreover, despite the vast amounts of data at the edge in general. The typical edge device generates insufficient data to facilitate the training of Machine Learning models from the ground up. To overcome these challenges and constraints, various techniques have been proposed such as Federated Learning (FL), Knowledge Distillation (KD), and Transfer Learning (TL), etc. These methods aim to optimize ML models to fit within the constraints of edge devices, thereby rendering them suitable for training in resource-constrained environments, scenarios with low data availability, or through collaborative training across multiple edge devices that leverage their collective data.
@@
This survey paper aims to provide a general overview on Edge Learning (EL), where we look into its methodologies, requirements, applications, challenges, and open research directions. We explore the state-of-the-art techniques and approaches used to train and optimize ML models in edge devices, highlighting their advantages and making a broad comparison between them. Additionally, we investigate the diverse range of applications that benefit from Edge Learning, and the frameworks, libraries and simulation tools used to optimize and enable edge learning.
@@
This survey is separated into eight different parts, excluding this Introduction and the Conclusions section at the end. First, Section II defines the terms edge learning and edge devices in more details, followed by the comparison of our survey with the existing ones highlighting our contributions in Section III, and the requirements and metrics for training ML models at the edge in Section IV. In Section (V), we explore the techniques that are used to enable, optimize and accelerate edge learning, and then make a detailed comparison between them in Section V-E. We explore how are the different types of ML techniques (Unsupervised learning, reinforcement learning, etc) are integrated with EL in Section VI, for the purpose of using these learning types in EL or to optimize or to enable the training of other models. In Section VII, we explore the use cases and current applications of edge learning. Then in Section VIII, we present different tools, frameworks and libraries used to train ML models at the edge (VIII-A) and create simulations and emulations for EL in Section VIII-B. Finally, in section IX, we highlight some open challenges, and try to predict future trends and research directions on edge learning.
@@ II Edge computing and Edge learning
In this section, we will introduce the concepts of Edge Computing, Edge Machine Learning and Edge Learning.
@@ II-A Edge Computing
Edge Computing (EC) is a new computing paradigm that aims to address the limitations of traditional Cloud Computing models in handling large scale data generated by the increasing number of smart devices connected to the Internet. It involves performing calculations at the edge of the network, closer to the user and the source of the data. EC emphasizes local, small-scale data storage and processing, providing benefits such as reduced bandwidth load, faster response speed, improved security, and enhanced privacy compared to traditional Cloud Computing models [2].
@@
  • Reduced latency: EC brings data processing closer to the source, reducing the time it takes for data to travel to a centralized cloud server, thereby reducing latency and improving response times [3].
  • Bandwidth optimization: EC reduces the need for transmitting large amounts of data to centralized cloud servers, resulting in reduced bandwidth load and reduced network congestion [2].
  • Improved data privacy: EC allows for local data processing, reducing the need to transmit sensitive data to centralized cloud servers, thereby minimizing the risk of data breaches and unauthorized access [2].
  • Operational resilience: EC enables applications to continue functioning even in disconnected or low-bandwidth environments, ensuring operational resilience and reducing dependency on a centralized cloud infrastructure [3].
The figure 1 shows the general architecture of EC, inspired by the ones proposed in [4] and [5]. We define edge devices as both edge servers and end devices, as well as other types of devices that weren’t specifically mentioned in the diagram, such as routers and routing switches.
@@ II-B Edge Learning
A promising development in the field of EC involves the integration of AI and ML. The concept of Edge Machine Learning (Edge ML) strives to facilitate the training and deployment of ML models directly on edge devices, encompassing both edge learning and edge inference. EL is characterized by the direct training of ML models on edge devices, thereby diminishing the reliance on centralized cloud infrastructure for model training. On the other hand, Edge Inference pertains to the deployment of ML models and the facilitation of inference on resource-constrained edge devices, regardless of whether the models were trained on the edge or within a cloud-based environment [6].
Most of the work done on Edge ML today involves edge inference [7, 8], however Edge Learning (EL) is still a promising paradigm, as it allows for localized model training, tailored to the specific requirements and resource constraints of edge devices, making it ideal for application requiring privacy preservation, and model tailoring for specific use cases.
Edge Learning employs various strategies, most of them are either categorized as distributed or collaborative learning methods that aim to distribute the training of ML model to multiple edge devices, such as federated or split learning; and on-device learning that aim to train a ML model on a single edge device, using either optimization or fine-tuning techniques if needed.
In this survey, we will delve into both on-device learning and distributed learning on edge devices. Distributed learning in EL is defined as the training of ML models collaboratively across multiple edge devices. While on-device learning, refers to the training of ML models in a single edge device. To ensure clarity, our definition of edge devices also encompasses edge servers, network elements, and end devices. We comprehensively address machine learning model training across all these devices.
III Comparison with existing surveys
There has been considerable surveys and reviews about Edge ML and Edge AI that attempts to define the field and present the different approaches that exist for AI in Edge. Most of them focus on edge inference or focus on a single aspect of edge learning, such as Federated Learning [9, 10] or on-device training [11, 12].
As defined in Section I, the purpose of this survey is to give a complete overview on the training ML models on the edge. This survey has multiple contributions that we will use to compare with the existing ones:
@@ III Comparison with existing surveys
    Explore techniques: Exploring the different techniques used to train or facilitate the training of ML models on edge devices.
@@
    Metrics for Edge Learning: Defining metrics to evaluate and compare EL approaches, and also highlight the requirements necessary to use edge learning in real life scenarios.
@@
    Compare techniques: Compare the different learning techniques previously presented based on the metrics and requirements, as well as the usage and popularity of the techniques in academia in the context of EL.
@@
    Explore Types of ML: Explore the different types of ML such as unsupervised or reinforcement learning in the context of EL.
@@
    Explore tools and libraries: Explore the tools and libraries used for training ML models on edge devices and providing simulations and emulators for EL.
@@
    Use-cases and applications: Present different use-cases and applications for edge learning explored in Academia.
The Table I present the relevant studies related to Edge ML, and compare them to our survey based on the aforementioned points.
TABLE I: Summary of Edge Machine Learning related surveys
Chen et al. [4]
2019
Wang et al. [13]
2020
Shi et al. [14]
2020
Leon Veas et al. [15]
2021
Dhar et al. [12]
2021
$\circ$
$\circ$
Tak et al. [16]
2021
$\circ$
Murshed et al. [17]
2022
Abreha et al. [9]
2022
$\circ$
$\circ$
$\circ$
Boobalan et al. [10]
2022
$\circ$
$\circ$
Joshi et al. [18]
2022
Cai et al. [19]
2022
Cui et al. [20]
2022
$\circ$
$\circ$
Imteaj et al. [21]
2022
$\circ$
$\circ$
$\circ$
Singh et al. [22]
2023
Li et al. [23]
2023
Hua et al. [24]
2023
Zhu et al. [11]
2023
$\circ$
$\circ$
Wu et al. [25]
2023
$\circ$
$\circ$
Our survey
2024
✓: Paper covers the topic fully for Edge Learning;
$\circ$ : Paper covers the topic partially for Edge Learning;
•: Paper covers the topic with a focus on both Edge Learning and Inference;
\: Paper covers the topic only from an Edge Inference point of view;
✗: Paper does not cover the topic;
IV Requirements for edge learning
The successful training of ML models at the edge requires meeting specific requirements that dictate the efficacy and performance of these models. These requirements can be viewed as metrics or evaluations, and they are essential for ensuring that the models perform optimally and efficiently within the resource-constrained environment inherent to edge devices. While there is no single metric to define the efficiency of training ML models in the edge [26], different ones can be used to make an estimation of how well the model will perform. We define some of these metrics/requirements in the figure 2.
Figure 2: Requirements of Edge Learning
  1. 1.
    Computational Efficiency: Computational Efficiency refers to the ability of an algorithm to achieve high performance with minimal computational cost. This is especially important in the context of EL, as edge devices, have limited computational resources [7], especially mobile edge devices which are even more resource-poor relative to static devices [27, 28]; and the training of ML models typically require high computational complexity [29].
  2. 2.
    Memory Footprint Efficiency: Similarly to computational complexity, edge devices often have low memory availability [30, 31], in contrast with the large memory footprint of ML models. This makes, memory efficiency an important requirement for EL.
  3. 3.
    Fast Training Time: Fast Training Time pertains to the rapid convergence of model parameters during the training phase. The importance of fast training time lies in its direct impact on the efficiency and responsiveness of edge devices. Edge devices, often characterized by limited computational capabilities, require ML models to be trained swiftly to minimize the processing burden and reduce energy consumption. Achieving Fast Training Time also ensures that models can be quickly adapted to evolving data patterns and ensures responsiveness and adaptivity where models can be efficiently updated to address dynamic environments and changing user requirements.
  4. 4.
    Optimized Bandwidth: Optimized Bandwidth refers to the optimization of the amount of data transferred in and out from edge devices, and the improvement of communication efficiency between edge devices. Bandwidth optimization is particularly important for distributed learning techniques at the edge and especially in bandwidth-limited systems, since these techniques require frequent sharing of the ML model across the network devices.
  5. 5.
    Low Energy Consumption: Energy consumption for edge devices is a crucial consideration, especially in mobile edge computing, considering the limited energy available on such devices. For this reason, ML models trained at the edge, need to be energy efficient, allowing for better computing performance and battery lifetime while ensuring the successful training of the models. Energy efficiency refers to the ability to use less energy to perform a specific task or function. It involves reducing energy waste and optimizing energy consumption.
  6. 6.
    Labelled Data Independency: The majority of the data produced at the edge are unlabelled data [32]. For that reason, it might be beneficial to use in EL, ML techniques that are fully or partially trained using unlabelled data such as unsupervised learning (Section VI-A), self-supervised learning (Section VI-D) or semi-supervised learning (Section VI-C).
@@ IV Requirements for edge learning
    Task Specific Metrics and Performance: As Edge learning encompass different ML tasks and use-cases. Specific metrics and benchmarks are used in general for the evaluation of the performances of the model and are vital for assessing its effectiveness in achieving its intended goals.
V Overview of Edge Learning techniques
Ultimately, the possibility of training an ML model on the edge depends on the resource requirements of the model, and the resources available on the device. With edge devices having increasingly more processing power, energy storage, and memory capacity available [33], it is currently possible to train a small ML model in an edge device without having to particularly optimize it for that regard or distribute the training across multiple devices. For example, KMeans in [34], Self-Organizing Map in [35] and SVM in [36]. However the training of more complex models that require heavier resources, such as neural networks is more challenging at the edge. Therefore, in this section, we will present an overview of techniques used to train more complex ML models in the edge, either through distributing the training to multiple devices, by training part of the model on the cloud and fine-tuning it on the local device or by optimizing or compressing the model to facilitate the training on the edge.
The Figure 3 shows a global view of Edge Learning techniques reviewed in this paper. The techniques are separated into four categories: (i) Distributed or collaborative techniques, such as federated or split learning; (ii) Techniques that rely on fine-tuning of a model trained on the cloud, such as Transfer or incremental learning; (iii) techniques that compress models to facilitate or support the training on the edge, such as quantization and knowledge distillation; (iiii) And Finally the other optimization te chniques that don’t fit needlessly into the previous categories.

### v3

@@ Abstract
Edge computing has gained significant traction in recent years, promising enhanced efficiency by integrating artificial intelligence capabilities at the edge. While the focus has primarily been on the deployment and inference of Machine Learning (ML) models at the edge, the training aspect remains less explored. This survey, explores the concept of edge learning, specifically the optimization of ML model training at the edge. The objective is to comprehensively explore diverse approaches and methodologies in edge learning, synthesize existing knowledge, identify challenges, and highlight future trends. Utilizing Scopus and Web of science advanced search, relevant literature on edge learning was identified, revealing a concentration of research efforts in distributed learning methods, particularly federated learning. This survey further provides a guideline for comparing techniques used to optimize ML for edge learning, along with an exploration of the different frameworks, libraries, and simulation tools available. In doing so, the paper contributes to a holistic understanding of the current landscape and future directions in the intersection of edge computing and machine learning, paving the way for informed comparisons between optimization methods and techniques designed for training on the edge.
@@
Machine Learning; Edge Computing; Edge AI; Edge Learning; On-Device Training; Edge intelligence; Artificial Intelligence; IoT.
publicationid: pubid: 0000–0000/00\$00.00 © 2024 IEEE
@@ I Introduction
In recent years, the fields of Artificial Intelligence (AI) and Machine Learning (ML) have witnessed significant growth, and have demonstrated remarkable success across various industrial applications [1]. ML’s essence lies in the interplay between algorithmic models and large quantities of data, as the latter is often required to successfully train ML models. Traditionally, datasets have been collected in cloud storage, databases, and data lakes. These datasets are then processed in central cloud servers to train various ML models.
@@
Conversely, the rapid proliferation of smart devices and sensors in recent years has led to an explosion of data generation at the edge of the network. With edge devices generating vast quantities of data closer to the source, growing concerns about privacy and security, as well as the desire to optimize the bandwidth consumption on the increasing number of edge devices and reduce the computational load on cloud servers, have driven a paradigm shift towards edge computing. In this context, computational processes are decentralized and migrated to edge devices. This sets the stage for a novel intersection between ML and edge computing.
@@
This shift has sparked growing interest towards edge ML. A union between machine learning and edge computing, deploying ML models at the edge, closer to end devices, enabling inference or training to occur at the edge. Edge learning is a subset of Edge ML that involves training ML models directly at the edge. Traditionally, ML models have relied on cloud infrastructure for training and deployment. However, this approach poses several challenges. These include high latency, significant communication overheads, and concerns around data privacy and security. By processing data closer to its source, edge learning tackles these challenges while enabling real-time decision-making and reducing cloud resource usage. Furthermore, this enables innovative ML applications, such as privacy-aware recommendation systems and smart technologies. These applications span multiple industries, including healthcare, manufacturing, agriculture, and space exploration.
@@
Training ML models at the edge poses unique challenges due to edge devices’ limited computational power and memory. Moreover, despite the abundance of data at the edge, individual devices usually lack sufficient data to train ML models from scratch. To address these challenges, techniques like federated learning, knowledge distillation, and transfer learning have been proposed. These methods aim to optimize ML models to fit within the constraints of edge devices, thereby rendering them suitable for training in resource-constrained environments, scenarios with low data availability, or through collaborative training across multiple edge devices that leverage their collective data.
@@
This survey paper aims to provide an overview on edge learning, covering its methodologies, requirements, applications, challenges, and open research directions. We explore state-of-the-art techniques for training and optimizing ML models on edge devices, highlighting their advantages. Furthermore, we compare these approaches, providing a broad overview of their strengths and weaknesses. We also examine the various applications that benefit from edge learning, as well as the frameworks, libraries, and simulation tools that support and optimize it. Please note that a background in ML and deep learning is assumed for this survey, and readers without this knowledge may find it helpful to consult a general introduction to the field, such as the ones found in [2, 3, 4, 5].
I-A Comparison with existing surveys
There have been considerable surveys about Edge ML that attempts to define the field and present the different approaches that exist for AI in Edge. Most of these surveys focus on edge inference or on a single aspect of edge learning, such as federated learning [6, 7] or on-device training [8, 9].
As previously defined, this survey provides a comprehensive overview of training ML models on edge devices. This survey has multiple contributions that we will use to compare with the existing surveys. The contributions are presented in the following five points that we will label as ”topics”.
  1. 1.
    Explore techniques: We examine various techniques for optimizing the training of ML models on edge devices.
  2. 2.
    Metrics for Edge Learning: We define metrics to evaluate and compare edge learning approaches, and identify requirements for edge learning in real-world scenarios.
  3. 3.
    Compare techniques: We compare the different edge learning techniques based on their performance, requirements, and popularity.
  4. 4.
    Explore Types of ML: We examine various types of ML, including unsupervised and reinforcement learning, in the context of edge learning.
  5. 5.
    Explore tools and libraries: We survey tools and libraries for training ML models on edge devices, as well as simulations and emulators for edge learning.
  6. 6.
    Use-cases and applications: We present various use cases and applications of edge learning explored in academic research.
Table I present the relevant studies related to Edge ML, and compare them to our survey based on the aforementioned topics. The symbols used in the table convey the extent to which each study addresses the different topics of edge learning outlined previously.
  1. 1.
    ✓ indicates that a survey comprehensively covers a particular topic in the context of edge learning.
  2. 2.
    $\circ$ denotes partial coverage of the topic, where a study may focus on a specific subset of the topic. For example, the surveys [10, 6, 7] explores techniques for using/optimizing ML for the edge but only focus on federated learning.
  3. 3.
    • signifies that a survey touches on the topic, but its primary emphasis lies on inference on the edge, rather than training, which often results in less comprehensive coverage of the training aspects.
  4. 4.
    ✗ indicates that a study does not address the topic at all.
TABLE I: Summary of Edge Machine Learning related surveys
Chen et al. [11] 2019
Wang et al. [12] 2020
Shi et al. [13] 2020
Xu et al. [14] 2020
Lim et al. [15] 2020 $\circ$
Leon Veas et al. [16] 2021
Dhar et al. [9] 2021 $\circ$ $\circ$
Tak et al. [10] 2021 $\circ$
Zhang et al. [17] 2021
Murshed et al. [18] 2022
Abreha et al. [6] 2022 $\circ$ $\circ$ $\circ$
Boobalan et al. [7] 2022 $\circ$ $\circ$
Joshi et al. [19] 2022
Cai et al. [20] 2022
Cui et al. [21] 2022 $\circ$ $\circ$
Imteaj et al. [22] 2022 $\circ$ $\circ$ $\circ$
Mendez et al. [23] 2022
Filho et al. [24] 2022
Ray et al. [25] 2022
Li et al. [26] 2023
Hua et al. [27] 2023
Zhu et al. [8] 2023 $\circ$ $\circ$
Wu et al. [28] 2023 $\circ$ $\circ$
Hoffpauir et al. [29] 2023
Barbuto et al. [30] 2023
Trinade et al. [31] 2024 $\circ$
Grzesik et al. [32] 2024
Jouini et al. [33] 2024
Our survey 2024
✓: Fully covers the topic of Edge Learning;
$\circ$ : Partially covers the topic of Edge Learning;
•: Focuses on both Edge Learning and Inference;
✗: Does not cover edge learning at all;
I-B Structure of the survey
@@
This survey is organized into six main sections, excluding this introduction and the conclusion (Section VIII). First, Section II provides a detailed definition of edge computing, edge learning and edge devices, and present the requirements and metrics for training ML models at the edge. In Section III, we explore the techniques used to enable, optimize, and accelerate edge learning. A detailed comparison between these techniques is presented in Section III-E. In Section IV, we discuss the integration of different types of ML such as unsupervised learning or reinforcement learning in the edge, to leverage these techniques in edge learning, optimize their performance, or enable the training of other models. In Section V, we explore the use cases and current applications of edge learning. Then in Section VI, we present different tools, frameworks and libraries used to create simulations and train ML models at the edge. Finally, Section VII identifies open challenges and discusses potential future trends and research directions in edge learning.
@@ II Edge computing and Edge learning
This section introduces the fundamental concepts of edge computing, edge machine learning, and edge learning. We will then examine the essential requirements of edge learning.
@@ II-A Edge Computing
Edge computing is a new computing paradigm that aims to address the limitations of traditional cloud computing models in handling large scale data generated by the increasing number of smart devices connected to the Internet. It involves performing calculations at the edge of the network, closer to the user and the source of the data. Edge computing emphasizes local, small-scale data storage and processing, providing benefits such as reduced bandwidth load, faster response speed, improved security, and enhanced privacy compared to traditional cloud computing models [34].
@@
  • Reduced latency: Edge computing brings data processing closer to the source, reducing the time it takes for data to travel to a centralized cloud server, thereby reducing latency and improving response time [35].
  • Bandwidth reduction: Edge computing reduces the need for transmitting large amounts of data to centralized cloud servers, resulting in reduced bandwidth load and reduced network congestion [34].
  • Improved data privacy: Edge computing allows for local data processing, reducing the need to transmit sensitive data to centralized cloud servers, thereby minimizing the risk of data breaches and unauthorized access [34].
  • Operational resilience: Edge computing enables applications to continue functioning even in disconnected or low-bandwidth environments, ensuring operational resilience and reducing dependency on centralized cloud infrastructures [35].
Figure 1 shows the general architecture of edge computing, inspired by the ones proposed in [11] and [36]. We define edge devices as both edge servers and end devices, as well as other types of devices that weren’t specifically mentioned in the diagram, such as routers and routing switches. For a deeper dive into edge computing, the reader can consult to edge computing surveys such as [37, 34].
@@ II-B Edge Learning
A key advancement in edge computing is the integration of AI and ML. Edge ML enables the training and deployment of ML models directly on edge devices, which includes both edge learning and edge inference. Edge learning, also known as edge training, involves training ML models directly on edge devices, reducing the reliance on centralized cloud infrastructure. In contrast, edge inference focuses on facilitating the inference of ML models on resource-constrained edge devices, regardless of where the models were trained [38]. Another term commonly used in the literature is edge intelligence, which shares similarities with edge ML. However, it also includes data collection, caching, processing, and analysis at the edge, making it a broader concept than edge ML [14].
While most Edge ML research focuses on edge inference [39, 40], edge learning remains a promising approach. By enabling localized model training, edge learning can be tailored to the specific requirements and resource constraints of edge devices, making it ideal for applications that require privacy preservation and model customization for specific use cases.
Edge Learning employs various strategies, most of them are either categorized as distributed or collaborative learning methods, which distribute the training of ML models across multiple edge devices, such as federated or split learning; and on-device learning, which involves training ML models on individual edge devices, and may employ optimization or fine-tuning techniques as necessary.
In this survey, we will explore both on-device learning and distributed learning on edge devices. Distributed learning is defined as the training of ML models collaboratively across multiple devices. In contrast, on-device learning refers to the training of ML models in a single device. To ensure clarity, our definition of edge devices also encompasses edge servers, network elements, and end devices. We comprehensively address ML model training across all these devices.
II-C Requirements for edge learning
The successful training of ML models at the edge requires meeting specific requirements that dictate the efficiency and performance of these models. These requirements are essential for ensuring that the models perform optimally and efficiently within the resource-constrained environment inherent to edge devices. While there is no single metric to define the efficiency of training ML models in the edge [41], different ones could be constructed be used to evaluate if and how well the aforementioned requirements are met, and estimate the performances of the model in resource’s contained environment.
@@ II-C Requirements for edge learning
    Computational Efficiency: Computational Efficiency refers to the ability of an algorithm to achieve high performance with minimal computational cost. This is especially important in the context of edge learning, as edge devices often have limited computational resources [39], and ML models typically require high computational complexity for their training [42].
@@
    Memory Footprint Efficiency: Similarly to computational complexity, edge devices often have low memory availability [43, 44], which contrast with the large memory requirements of ML models.
@@
    Fast Training Time: Fast training time refers to the rapid convergence of model parameters during the training phase. Fast training time is crucial for edge devices, as it directly impacts their efficiency and responsiveness. Edge devices are often characterized by limited computational capabilities, as mentioned earlier. As such, they require ML models to be trained swiftly to minimize the processing burden and reduce energy consumption. Fast training time also enables models to adapt quickly to changing data patterns, ensuring responsiveness and adaptivity. This allows models to be efficiently updated to address dynamic environments and changing user requirements.
@@
    Minimized Bandwidth Consumption: Reducing bandwidth usage involves minimizing data transfer between edge devices and improving communication efficiency among them. This is particularly important for distributed learning techniques and especially in bandwidth-limited systems, since these techniques require frequent sharing of the ML model across the network devices.
@@
    Low Energy Consumption: Energy consumption is a crucial consideration for edge devices, especially in mobile edge computing. This is due to the limited energy available on such devices. Therefore, ML models trained at the edge must be energy-efficient to ensure better computing performance, longer battery life, and successful model training. Energy efficiency refers to the ability to perform tasks or functions using minimal energy. It involves reducing energy waste and optimizing energy consumption.
@@
    Labelled Data Independency: Most edge-generated data is unlabelled [45]. Therefore, using ML techniques that can handle unlabeled data, such as unsupervised (Section IV-A), self-supervised (Section IV-D), or semi-supervised learning (Section IV-C), may be beneficial in edge learning.
@@
    Task Specific Metrics and Performance: Since edge learning encompass different ML tasks and use-cases. Specific metrics and benchmarks are commonly used to evaluate a model’s performance to assess its effectiveness in achieving its intended goals.
III Overview of Edge Learning techniques
In general, edge ML training is similar to traditional ML training, with the added requirements and constraints outlined in Section II-C. The feasibility of training on the edge depends on the resource requirements of the model, and the device resources. Today, the increasing processing power, energy storage, and memory capacity of edge devices [46] enables small ML models to be trained on edge devices without requiring significant optimization or distribution. For example, KMeans in [47], Self-Organizing Map in [48] and SVM in [49]. However, the training of more complex models that require heavier resources, such as neural networks, is more challenging at the edge. Therefore, in this section, we will present an overview of techniques used to optimize the training of more complex ML models on the edge. These techniques include distributing training across multiple devices, cloud-based training with local fine-tuning, and model optimization or compression to enable edge training.
Figure 2 shows a global view of edge learning techniques reviewed in this paper. The techniques are separated into four categories: (i) Distributed or collaborative techniques, such as federated or split learning; (ii) Techniques that rely on fine-tuning of a model trained on the cloud, such as Transfer or incremental learning; (iii) techniques that compress models to facilitate or support the training on the edge, such as quantization and knowledge distillation; (iv) And finally the other optimization techniques that don’t fit neatly into the previous categories.